OpenStudy (anonymous):

Check my answer? Write the equation of the line passing through point (2,-3) and is perpendicular to y=4x+7 in standard form. I got 4x+y=-11; is it right?

hartnn (hartnn):

no again...

hartnn (hartnn):

slope of given line ?

OpenStudy (anonymous):

4

hartnn (hartnn):

yes, so slope of perpendicular line ?

OpenStudy (anonymous):

4/1 ?

hartnn (hartnn):

the product of slope of perpendicular lines = -1

hartnn (hartnn):

so m1 * m2 = -1 m1 = 4, m2= ?

OpenStudy (anonymous):

oh 4!

OpenStudy (anonymous):

-4 my bad

hartnn (hartnn):

no, slope of perpendicular line will be negative *reciprocal*

hartnn (hartnn):

so m1*m2 = -1 4m2=-1 m2 = -1/4 ok ?

OpenStudy (anonymous):

ohk sorry I'm not really good at this I just learned this

hartnn (hartnn):

so now u have slope of required line and a point on it, which equation will u use ?

OpenStudy (anonymous):

point-slope form? y-y1=m(x-x1)

hartnn (hartnn):

thats correct! so what u get ?

OpenStudy (anonymous):

y-(-3)=-1/4(x-2) ?

hartnn (hartnn):

correct :) now simplify....

OpenStudy (anonymous):

y-(-3)=-1/4x - idk/: 0.5?

hartnn (hartnn):

(-1/4)(-2) = +2/4 = +0.5

hartnn (hartnn):

y-(-3) = y+3

OpenStudy (anonymous):

oh wait.. y+3 = -1/4x - 0.5 then would you add -1/4x to the left side? since I have to write it in standard form? ax+by=c?

hartnn (hartnn):

1) its y+3 = -1/4x + 0.5 2) yes., add 1/4 x , not -1/4 x

OpenStudy (anonymous):

ohk so .. 1/4x+y+3=0.5 then add subtract 3 on both sides?

hartnn (hartnn):

yes, correct

OpenStudy (anonymous):

1/4x+y=2.5

hartnn (hartnn):

0.5-3 = -2.5

hartnn (hartnn):

or u could first multiply by 4 on both sides

OpenStudy (anonymous):

oh yeah.. sorry.. and oh ohk if i did that it would be x+y=2

hartnn (hartnn):

let me show you steps: y+3 = -1/4x + 0.5 y+1/4 x +3= 0.5 multiplying 4 on both sides 4y+ x + 12 = 2 now can u solve ahead

OpenStudy (anonymous):

4y+x=14

hartnn (hartnn):

2-12=?

OpenStudy (anonymous):

oh whoops -10

hartnn (hartnn):

you will subtract 12 from both sides, right ?

OpenStudy (anonymous):

yes, sorry

hartnn (hartnn):

so finally, 4y+x=-10

OpenStudy (anonymous):

yeah thank you though! (:

hartnn (hartnn):

welcome ^_^